A standardized​ exam's scores are normally distributed. In a recent​ year, the mean test score was 20.5 and the standard deviation was 5.4. The test scores of four students selected at random are 15​, 23​, 8​, and 34. Find the​ z-scores that correspond to each value and determine whether any of the values are unusual.

The​ z-score for 15 is

The z- score for 23 is

The z-score for 8 is

The z-score for 34 is

Which​ values, if​ any, are​ unusual? Select the correct choice below​ and, if​ necessary, fill in the answer box within your choice.
Answer #1

Solution :

z = (x - ) /

The​ z-score for 15 is = (15 - 20.5) / 5.4 = -1.02

The z- score for 23 is = (23 - 20.5) / 5.4 = 0.46

The z-score for 8 is = (8 - 20.5) / 5.4 = -2.31

The z-score for 34 is = (34 - 20.5) / 5.4 = 2.5

The​ z-score for 15 and the z-score for 34 is are unusual .